OpenStudy (anonymous):

You draw 2 cards from a well-shuffled deck of 52 cards. After you draw the first cards, you do not replace it in the deck. That is the probability of P(get ace of spades on first draw and get queen of hearts on second draw)?

OpenStudy (mary.rojas):

@phi help? :)

OpenStudy (mary.rojas):

@goformit100 help?

OpenStudy (mary.rojas):

@campbell_st help?

OpenStudy (campbell_st):

so isn't it P(ace spades) 1/52 P(Q or Hearts) 1/51 so no replacement after 1st draw P(Ace + Q) = 1/52 * 1/51

OpenStudy (anonymous):

so 1/2652?

OpenStudy (campbell_st):

well thats provided the 1st drawn card isn't replaced.,

OpenStudy (anonymous):

ok so its right then?

OpenStudy (campbell_st):

yep... that would be my answer

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

There is a 40 percent chance of rain today and a 50 percent chance of rain tomorrow. If the two events are independent, what is the probability that it will rain today or tomorrow? A. 60 B. 90 C. 72 D. 70 E. 18

OpenStudy (anonymous):

@mary.rojas

OpenStudy (campbell_st):

40% = 2/5 50% = 1/2 so its 2/5 x 1/2 and then change it to a percentage

OpenStudy (mary.rojas):

P(rains today or rains tomorrow) = P(rains today, rains tomorrow) + P(rains today, doesn't rain tomorrow) + P(doesn't rain today, rains tomorrow) = (2/5 * 1/2) + (2/5 * 1/2) + (3/5 * 1/2) = 2/10 + 2/10 + 3/10 = 7/10 = 70% So, there is a 70% chance that it will rain today or tomorrow (or both). So, your answer is D. Hope this helps! :)

OpenStudy (anonymous):

ok.. 70% it is then

OpenStudy (mary.rojas):

wait, is it right?

OpenStudy (anonymous):

i think i ll tell u if its right. im gonna post a new question though. ill tag u